OpenStudy (anonymous):

What is the inverse of h? h(x) = 6x + 1

OpenStudy (anonymous):

could have sworn i just answered this one \[x=6y+1\] \[x-1=6y\] \[y=\frac{x-1}{6}\]

OpenStudy (anonymous):

y=6x+1 x=6y+1 -1 -1 x-1=6y y=x-1/6\[\frac{ x-1 }{ 6y }\]

OpenStudy (anonymous):

this is not one of the multiple choice ancers